# DualNILM: Energy Injection Identification enabled Disaggregation with Deep Multi-Task Learning

### v2

@@ E-B Laboratory Dataset Results
Accuracy (%) | Recall (%) | Precision (%) | F1 (%) | ||
| Air compressor | FHMM | 99.56 | 14.29 | 14.29 | 14.29 | ||
| XGBoost | 99.55 | 0.00 | 0.00 | 0.00 | |||
| Seq2Point | 99.52 | 0.00 | 0.00 | 0.00 | |||
| CNN-LSTM | 99.52 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 99.52 | 0.00 | 0.00 | 0.00 | |||
| UNetNILM | 99.52 | 0.00 | 0.00 | 0.00 | |||
| DualNILM | 99.97 | 83.77 | 83.40 | 83.52 | |||
| Air conditioner | FHMM | 60.46 | 0.06 | 14.29 | 0.11 | ||
| XGBoost | 85.28 | 79.10 | 78.25 | 77.24 | |||
| Seq2Point | 41.63 | 100.00 | 41.63 | 55.46 | |||
| CNN-LSTM | 63.85 | 99.40 | 52.71 | 66.29 | |||
| Transformer | 73.70 | 99.92 | 61.77 | 73.58 | |||
| UNetNILM | 84.72 | 68.70 | 71.17 | 69.50 | |||
| DualNILM | 99.57 | 99.41 | 98.94 | 99.17 | |||
| Air purifier | FHMM | 36.96 | 48.72 | 36.41 | 36.37 | ||
| XGBoost | 58.30 | 44.27 | 45.73 | 40.50 | |||
| Seq2Point | 42.71 | 85.71 | 42.71 | 53.99 | |||
| CNN-LSTM | 42.71 | 85.71 | 42.71 | 53.99 | |||
| Transformer | 42.71 | 85.71 | 42.71 | 53.99 | |||
| UNetNILM | 74.15 | 60.96 | 66.36 | 60.32 | |||
| DualNILM | 93.58 | 81.38 | 80.01 | 80.38 | |||
| Heater | FHMM | 54.31 | 85.71 | 37.34 | 48.05 | ||
| XGBoost | 95.37 | 81.65 | 80.21 | 80.51 | |||
| Seq2Point | 31.97 | 85.71 | 31.97 | 43.00 | |||
| CNN-LSTM | 54.57 | 85.65 | 37.94 | 48.92 | |||
| Transformer | 99.01 | 84.77 | 83.94 | 84.29 | |||
| UNetNILM | 98.33 | 82.59 | 85.00 | 83.62 | |||
| DualNILM | 98.77 | 84.88 | 79.08 | 81.11 | |||
| Light bulb Type I | FHMM | 74.81 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 87.75 | 4.26 | 23.20 | 6.88 | |||
| Seq2Point | 88.53 | 0.00 | 0.00 | 0.00 | |||
| CNN-LSTM | 88.53 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 88.53 | 0.00 | 0.00 | 0.00 | |||
| UNetNILM | 89.83 | 16.59 | 51.65 | 17.93 | |||
| DualNILM | 99.63 | 67.66 | 64.40 | 65.59 | |||
| Light bulb Type II | FHMM | 74.81 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 91.48 | 39.80 | 48.57 | 41.06 | |||
| Seq2Point | 90.35 | 0.00 | 0.00 | 0.00 | |||
| CNN-LSTM | 90.35 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 90.35 | 0.00 | 0.00 | 0.00 | |||
| UNetNILM | 94.97 | 50.57 | 68.62 | 56.37 | |||
| DualNILM | 99.05 | 81.57 | 83.89 | 82.63 | |||
| Micro-inverter | FHMM | 25.73 | 49.49 | 15.59 | 21.93 | ||
| XGBoost | 83.04 | 88.33 | 88.75 | 88.09 | |||
| Seq2Point | 78.68 | 100.00 | 78.68 | 87.46 | |||
| CNN-LSTM | 78.68 | 100.00 | 78.68 | 87.46 | |||
| Transformer | 78.68 | 100.00 | 78.68 | 87.46 | |||
| UNetNILM | 82.85 | 98.64 | 82.87 | 89.48 | |||
| DualNILM | 99.17 | 99.37 | 99.42 | 99.39 | |||
| Avg. | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | ||
| XGBoost | 85.82 | 48.20 | 52.10 | 47.75 | |||
| Seq2Point | 67.63 | 53.06 | 27.86 | 34.27 | |||
| CNN-LSTM | 74.03 | 52.97 | 30.29 | 36.67 | |||
| Transformer | 81.79 | 52.91 | 38.16 | 42.76 | |||
| UNetNILM | 89.20 | 54.01 | 60.81 | 53.89 | |||
| DualNILM | 98.54 | 85.43 | 84.16 | 84.54 | |||
| Injection Disaggregation | FHMM | Avg. RMSE: 0.5015 | Avg. MAE: 0.4492 | ||||
| Seq2Seq | Avg. RMSE: 0.3512 | Avg. MAE: 0.2622 | |||||
| DAE | Avg. RMSE: 0.8830 | Avg. MAE: 0.7851 | |||||
| UNetNILM | Avg. RMSE: 0.2305 | Avg. MAE: 0.1883 | |||||
| DualNILM | Avg. RMSE: 0.1429 | Avg. MAE: 0.0804 | |||||
E-C REDD Dataset Results
E-C1 Full Results on House 1
See Table X.
Accuracy (%) | Recall (%) | Precision (%) | F1 (%) | ||
| Microwave | FHMM | 99.13 | 0.06 | 33.33 | 0.12 | ||
| XGBoost | 99.40 | 44.51 | 72.75 | 52.40 | |||
| Seq2Point | 99.37 | 35.85 | 98.67 | 51.25 | |||
| CNN-LSTM | 99.11 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 99.11 | 0.00 | 0.00 | 0.00 | |||
| UNetNILM | 99.22 | 20.76 | 84.45 | 31.96 | |||
| DualNILM | 99.90 | 88.10 | 95.87 | 91.75 | |||
| Fridge | FHMM | 99.96 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 91.91 | 59.20 | 91.80 | 71.92 | |||
| Seq2Point | 93.99 | 80.91 | 84.44 | 82.58 | |||
| CNN-LSTM | 97.02 | 91.32 | 91.93 | 91.61 | |||
| Transformer | 84.18 | 78.06 | 53.95 | 63.72 | |||
| UNetNILM | 97.52 | 89.83 | 95.82 | 92.73 | |||
| DualNILM | 99.63 | 99.49 | 98.46 | 98.97 | |||
| Dish Washer | FHMM | 98.64 | 4.80 | 32.49 | 8.37 | ||
| XGBoost | 99.61 | 89.32 | 82.77 | 85.03 | |||
| Seq2Point | 98.75 | 0.00 | 0.00 | 0.00 | |||
| CNN-LSTM | 98.93 | 40.75 | 62.29 | 48.66 | |||
| Transformer | 99.12 | 76.22 | 65.42 | 69.89 | |||
| UNetNILM | 99.00 | 96.16 | 56.51 | 71.02 | |||
| DualNILM | 99.96 | 99.47 | 97.33 | 98.39 | |||
| Washing Machine | FHMM | 98.78 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 99.74 | 81.06 | 97.88 | 88.10 | |||
| Seq2Point | 98.67 | 0.00 | 0.00 | 0.00 | |||
| CNN-LSTM | 98.67 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 98.79 | 26.20 | 54.27 | 35.11 | |||
| UNetNILM | 99.54 | 73.82 | 89.40 | 80.68 | |||
| DualNILM | 99.85 | 94.26 | 94.23 | 94.23 | |||
| Micro-inverter | FHMM | 1.10 | 2.45 | 1.52 | 1.86 | ||
| XGBoost | 98.83 | 98.41 | 98.51 | 98.45 | |||
| Seq2Point | 97.46 | 94.22 | 99.01 | 96.55 | |||
| CNN-LSTM | 99.63 | 99.54 | 99.53 | 99.54 | |||
| Transformer | 98.78 | 99.93 | 96.86 | 98.36 | |||
| UNetNILM | 99.96 | 99.97 | 99.92 | 99.95 | |||
| DualNILM | 99.87 | 100.00 | 99.68 | 99.84 | |||
| Avg. | FHMM | 79.52 | 1.46 | 13.47 | 2.07 | ||
| XGBoost | 97.90 | 74.50 | 88.74 | 79.18 | |||
| Seq2Point | 97.65 | 42.20 | 56.42 | 46.07 | |||
| CNN-LSTM | 98.67 | 46.32 | 50.75 | 47.96 | |||
| Transformer | 95.99 | 56.08 | 54.10 | 53.42 | |||
| UNetNILM | 99.05 | 76.11 | 85.22 | 75.27 | |||
| DualNILM | 99.84 | 96.26 | 97.11 | 96.64 | |||
| Injection Disaggregation | FHMM | Avg. RMSE: 0.8093 | Avg. MAE: 0.2200 | ||||
| Seq2Seq | Avg. RMSE: 0.6834 | Avg. MAE: 0.2370 | |||||
| DAE | Avg. RMSE: 0.8115 | Avg. MAE: 0.2171 | |||||
| UNetNILM | Avg. RMSE: 0.7589 | Avg. MAE: 0.1983 | |||||
| DualNILM | Avg. RMSE: 0.6661 | Avg. MAE: 0.1777 | |||||
E-C2 Full Results on House 2
See Table XI.
Accuracy (%) | Recall (%) | Precision (%) | F1 (%) | ||
| Microwave | FHMM | 99.56 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 99.93 | 80.21 | 99.08 | 87.95 | |||
| Seq2Point | 99.51 | 91.55 | 42.30 | 57.73 | |||
| CNN-LSTM | 99.63 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 99.63 | 0.00 | 0.00 | 0.00 | |||
| UNetNILM | 99.70 | 24.67 | 77.10 | 37.26 | |||
| DualNILM | 99.91 | 79.18 | 93.96 | 85.67 | |||
| Fridge | FHMM | 99.99 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 98.97 | 98.34 | 99.35 | 98.84 | |||
| Seq2Point | 87.57 | 74.39 | 97.04 | 84.13 | |||
| CNN-LSTM | 97.34 | 97.27 | 96.73 | 97.00 | |||
| Transformer | 93.64 | 91.38 | 93.99 | 92.65 | |||
| UNetNILM | 99.05 | 99.32 | 98.53 | 98.92 | |||
| DualNILM | 99.15 | 99.94 | 98.20 | 99.06 | |||
| Dish Washer | FHMM | 99.30 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 99.92 | 99.48 | 90.22 | 94.61 | |||
| Seq2Point | 99.29 | 0.00 | 0.00 | 0.00 | |||
| CNN-LSTM | 99.29 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 99.29 | 0.00 | 0.00 | 0.00 | |||
| UNetNILM | 99.57 | 32.67 | 33.33 | 33.00 | |||
| DualNILM | 99.92 | 96.73 | 94.06 | 95.13 | |||
| Micro-inverter | FHMM | 1.46 | 2.31 | 1.36 | 1.71 | ||
| XGBoost | 94.91 | 91.83 | 93.68 | 92.73 | |||
| Seq2Point | 79.33 | 45.65 | 94.42 | 61.54 | |||
| Transformer | 99.13 | 99.73 | 97.92 | 98.82 | |||
| UNetNILM | 99.96 | 99.99 | 99.90 | 99.95 | |||
| DualNILM | 97.61 | 95.32 | 98.02 | 96.65 | |||
| Avg. | FHMM | 75.08 | 0.58 | 0.34 | 0.43 | ||
| XGBoost | 98.43 | 92.47 | 95.58 | 93.53 | |||
| Seq2Point | 91.43 | 52.90 | 58.44 | 50.85 | |||
| CNN-LSTM | 98.86 | 49.19 | 48.74 | 48.96 | |||
| Transformer | 97.92 | 47.78 | 47.98 | 47.87 | |||
| UNetNILM | 99.57 | 64.16 | 77.22 | 67.28 | |||
| DualNILM | 99.70 | 92.48 | 94.00 | 93.11 | |||
| Injection Disaggregation | FHMM | Avg. RMSE: 0.8403 | Avg. MAE: 0.3367 | ||||
| Seq2Seq | Avg. RMSE: 0.7217 | Avg. MAE: 0.3287 | |||||
| DAE | Avg. RMSE: 0.8448 | Avg. MAE: 0.3356 | |||||
| UNetNILM | Avg. RMSE: 2.2077 | Avg. MAE: 0.6601 | |||||
| DualNILM | Avg. RMSE: 0.7065 | Avg. MAE: 0.2768 | |||||
E-C3 Full Results on House 3
See Table XII.
Accuracy (%) | Recall (%) | Precision (%) | F1 (%) | ||
| Microwave | FHMM | 97.95 | 16.22 | 2.77 | 4.73 | ||
| XGBoost | 99.78 | 42.67 | 52.10 | 46.80 | |||
| Seq2Point | 99.80 | 25.54 | 66.67 | 34.37 | |||
| CNN-LSTM | 99.75 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 99.75 | 0.00 | 0.00 | 0.00 | |||
| UNetNILM | 99.70 | 0.00 | 0.00 | 0.00 | |||
| DualNILM | 99.88 | 65.70 | 93.19 | 75.20 | |||
| Fridge | FHMM | 97.33 | 4.74 | 1.44 | 2.21 | ||
| XGBoost | 87.11 | 67.22 | 97.50 | 79.56 | |||
| Seq2Point | 93.89 | 93.92 | 90.19 | 91.97 | |||
| CNN-LSTM | 85.64 | 68.17 | 93.58 | 78.25 | |||
| Transformer | 93.35 | 96.98 | 85.76 | 91.01 | |||
| UNetNILM | 96.51 | 99.66 | 91.25 | 95.26 | |||
| DualNILM | 98.82 | 99.00 | 98.13 | 98.56 | |||
| Dish Washer | FHMM | 97.65 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 99.81 | 24.60 | 32.31 | 27.93 | |||
| Seq2Point | 99.60 | 0.00 | 0.00 | 0.00 | |||
| CNN-LSTM | 99.60 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 98.96 | 11.46 | 10.17 | 10.78 | |||
| UNetNILM | 99.60 | 0.00 | 0.00 | 0.00 | |||
| DualNILM | 99.93 | 29.00 | 31.84 | 30.35 | |||
| Washing Machine | FHMM | 98.02 | 0.00 | 0.00 | 0.00 | ||
| XGBoost | 99.93 | 96.63 | 100.00 | 98.28 | |||
| Seq2Point | 98.18 | 0.00 | 0.00 | 0.00 | |||
| CNN-LSTM | 98.18 | 0.00 | 0.00 | 0.00 | |||
| Transformer | 98.18 | 0.00 | 0.00 | 0.00 | |||
| UNetNILM | 99.56 | 100.00 | 82.45 | 90.29 | |||
| DualNILM | 99.86 | 93.51 | 99.29 | 96.25 | |||
| Micro-inverter | FHMM | 28.27 | 1.67 | 0.30 | 0.50 | ||
| XGBoost | 87.80 | 0.00 | 0.00 | 0.00 | |||
| Seq2Point | 99.02 | 94.45 | 97.66 | 96.02 | |||
| CNN-LSTM | 96.40 | 76.06 | 97.27 | 84.87 | |||
| Transformer | 98.79 | 92.62 | 98.55 | 95.38 | |||
| UNetNILM | 99.99 | 99.90 | 99.96 | 99.93 | |||
| DualNILM | 99.96 | 99.90 | 99.83 | 99.86 | |||
| Avg. | FHMM | 83.84 | 4.53 | 0.90 | 1.49 | ||
| XGBoost | 94.89 | 46.22 | 56.38 | 50.51 | |||
| Seq2Point | 98.10 | 42.78 | 50.90 | 44.47 | |||
| CNN-LSTM | 95.92 | 28.85 | 38.17 | 32.62 | |||
| Transformer | 97.81 | 40.21 | 38.89 | 39.43 | |||
| UNetNILM | 99.07 | 59.91 | 54.73 | 57.10 | |||
| DualNILM | 99.69 | 77.42 | 84.46 | 80.04 | |||
| Injection Disaggregation | FHMM | Avg. RMSE: 0.5575 | Avg. MAE: 0.1766 | ||||
| Seq2Seq | Avg. RMSE: 0.4974 | Avg. MAE: 0.1940 | |||||
| DAE | Avg. RMSE: 0.5606 | Avg. MAE: 0.1743 | |||||
[…]

### v3

@@ E.2. Laboratory Dataset Results
| Appliance | Method | Accuracy (%) | Recall (%) | Precision (%) | F1 (%) |
| Air compressor | FHMM | 99.56 | 14.29 | 14.29 | 14.29 |
| XGBoost | 99.55 | 0.00 | 0.00 | 0.00 | |
| Seq2Point | 99.52 | 0.00 | 0.00 | 0.00 | |
| CNN-LSTM | 99.52 | 0.00 | 0.00 | 0.00 | |
| Transformer | 99.52 | 0.00 | 0.00 | 0.00 | |
| UNetNILM | 99.52 | 0.00 | 0.00 | 0.00 | |
| DualNILM | 99.97 | 83.77 | 83.40 | 83.52 | |
| Air conditioner | FHMM | 60.46 | 0.06 | 14.29 | 0.11 |
| XGBoost | 85.28 | 79.10 | 78.25 | 77.24 | |
| Seq2Point | 41.63 | 100.00 | 41.63 | 55.46 | |
| CNN-LSTM | 63.85 | 99.40 | 52.71 | 66.29 | |
| Transformer | 73.70 | 99.92 | 61.77 | 73.58 | |
| UNetNILM | 84.72 | 68.70 | 71.17 | 69.50 | |
| DualNILM | 99.57 | 99.41 | 98.94 | 99.17 | |
| Air purifier | FHMM | 36.96 | 48.72 | 36.41 | 36.37 |
| XGBoost | 58.30 | 44.27 | 45.73 | 40.50 | |
| Seq2Point | 42.71 | 85.71 | 42.71 | 53.99 | |
| CNN-LSTM | 42.71 | 85.71 | 42.71 | 53.99 | |
| Transformer | 42.71 | 85.71 | 42.71 | 53.99 | |
| UNetNILM | 74.15 | 60.96 | 66.36 | 60.32 | |
| DualNILM | 93.58 | 81.38 | 80.01 | 80.38 | |
| Heater | FHMM | 54.31 | 85.71 | 37.34 | 48.05 |
| XGBoost | 95.37 | 81.65 | 80.21 | 80.51 | |
| Seq2Point | 31.97 | 85.71 | 31.97 | 43.00 | |
| CNN-LSTM | 54.57 | 85.65 | 37.94 | 48.92 | |
| Transformer | 99.01 | 84.77 | 83.94 | 84.29 | |
| UNetNILM | 98.33 | 82.59 | 85.00 | 83.62 | |
| DualNILM | 98.77 | 84.88 | 79.08 | 81.11 | |
| Light bulb Type I | FHMM | 74.81 | 0.00 | 0.00 | 0.00 |
| XGBoost | 87.75 | 4.26 | 23.20 | 6.88 | |
| Seq2Point | 88.53 | 0.00 | 0.00 | 0.00 | |
| CNN-LSTM | 88.53 | 0.00 | 0.00 | 0.00 | |
| Transformer | 88.53 | 0.00 | 0.00 | 0.00 | |
| UNetNILM | 89.83 | 16.59 | 51.65 | 17.93 | |
| DualNILM | 99.63 | 67.66 | 64.40 | 65.59 | |
| Light bulb Type II | FHMM | 74.81 | 0.00 | 0.00 | 0.00 |
| XGBoost | 91.48 | 39.80 | 48.57 | 41.06 | |
| Seq2Point | 90.35 | 0.00 | 0.00 | 0.00 | |
| CNN-LSTM | 90.35 | 0.00 | 0.00 | 0.00 | |
| Transformer | 90.35 | 0.00 | 0.00 | 0.00 | |
| UNetNILM | 94.97 | 50.57 | 68.62 | 56.37 | |
| DualNILM | 99.05 | 81.57 | 83.89 | 82.63 | |
| Micro-inverter | FHMM | 25.73 | 49.49 | 15.59 | 21.93 |
| XGBoost | 83.04 | 88.33 | 88.75 | 88.09 | |
| Seq2Point | 78.68 | 100.00 | 78.68 | 87.46 | |
| CNN-LSTM | 78.68 | 100.00 | 78.68 | 87.46 | |
| Transformer | 78.68 | 100.00 | 78.68 | 87.46 | |
| UNetNILM | 82.85 | 98.64 | 82.87 | 89.48 | |
| DualNILM | 99.17 | 99.37 | 99.42 | 99.39 | |
| Avg. | FHMM | 60.95 | 28.32 | 16.84 | 17.25 |
| XGBoost | 85.82 | 48.20 | 52.10 | 47.75 | |
| Seq2Point | 67.63 | 53.06 | 27.86 | 34.27 | |
| CNN-LSTM | 74.03 | 52.97 | 30.29 | 36.67 | |
| Transformer | 81.79 | 52.91 | 38.16 | 42.76 | |
| UNetNILM | 89.20 | 54.01 | 60.81 | 53.89 | |
| DualNILM | 98.54 | 85.43 | 84.16 | 84.54 | |
| Injection Disaggregation | FHMM | Avg. RMSE: 0.5015 | Avg. MAE: 0.4492 | ||
| Seq2Seq | Avg. RMSE: 0.3512 | Avg. MAE: 0.2622 | |||
| DAE | Avg. RMSE: 0.8830 | Avg. MAE: 0.7851 | |||
| UNetNILM | Avg. RMSE: 0.2305 | Avg. MAE: 0.1883 | |||
| DualNILM | Avg. RMSE: 0.1429 | Avg. MAE: 0.0804 | |||
E.3. REDD Dataset Results
E.3.1. Full Results on House 1
See Table 9.
| Appliance | Method | Accuracy (%) | Recall (%) | Precision (%) | F1 (%) |
| Microwave | FHMM | 99.13 | 0.06 | 33.33 | 0.12 |
| XGBoost | 99.40 | 44.51 | 72.75 | 52.40 | |
| Seq2Point | 99.37 | 35.85 | 98.67 | 51.25 | |
| CNN-LSTM | 99.11 | 0.00 | 0.00 | 0.00 | |
| Transformer | 99.11 | 0.00 | 0.00 | 0.00 | |
| UNetNILM | 99.22 | 20.76 | 84.45 | 31.96 | |
| DualNILM | 99.90 | 88.10 | 95.87 | 91.75 | |
| Fridge | FHMM | 99.96 | 0.00 | 0.00 | 0.00 |
| XGBoost | 91.91 | 59.20 | 91.80 | 71.92 | |
| Seq2Point | 93.99 | 80.91 | 84.44 | 82.58 | |
| CNN-LSTM | 97.02 | 91.32 | 91.93 | 91.61 | |
| Transformer | 84.18 | 78.06 | 53.95 | 63.72 | |
| UNetNILM | 97.52 | 89.83 | 95.82 | 92.73 | |
| DualNILM | 99.63 | 99.49 | 98.46 | 98.97 | |
| Dish Washer | FHMM | 98.64 | 4.80 | 32.49 | 8.37 |
| XGBoost | 99.61 | 89.32 | 82.77 | 85.03 | |
| Seq2Point | 98.75 | 0.00 | 0.00 | 0.00 | |
| CNN-LSTM | 98.93 | 40.75 | 62.29 | 48.66 | |
| Transformer | 99.12 | 76.22 | 65.42 | 69.89 | |
| UNetNILM | 99.00 | 96.16 | 56.51 | 71.02 | |
| DualNILM | 99.96 | 99.47 | 97.33 | 98.39 | |
| Washing Machine | FHMM | 98.78 | 0.00 | 0.00 | 0.00 |
| XGBoost | 99.74 | 81.06 | 97.88 | 88.10 | |
| Seq2Point | 98.67 | 0.00 | 0.00 | 0.00 | |
| CNN-LSTM | 98.67 | 0.00 | 0.00 | 0.00 | |
| Transformer | 98.79 | 26.20 | 54.27 | 35.11 | |
| UNetNILM | 99.54 | 73.82 | 89.40 | 80.68 | |
| DualNILM | 99.85 | 94.26 | 94.23 | 94.23 | |
| Micro-inverter | FHMM | 1.10 | 2.45 | 1.52 | 1.86 |
| XGBoost | 98.83 | 98.41 | 98.51 | 98.45 | |
| Seq2Point | 97.46 | 94.22 | 99.01 | 96.55 | |
| CNN-LSTM | 99.63 | 99.54 | 99.53 | 99.54 | |
| Transformer | 98.78 | 99.93 | 96.86 | 98.36 | |
| UNetNILM | 99.96 | 99.97 | 99.92 | 99.95 | |
| DualNILM | 99.87 | 100.00 | 99.68 | 99.84 | |
| Avg. | FHMM | 79.52 | 1.46 | 13.47 | 2.07 |
| XGBoost | 97.90 | 74.50 | 88.74 | 79.18 | |
| Seq2Point | 97.65 | 42.20 | 56.42 | 46.07 | |
| CNN-LSTM | 98.67 | 46.32 | 50.75 | 47.96 | |
| Transformer | 95.99 | 56.08 | 54.10 | 53.42 | |
| UNetNILM | 99.05 | 76.11 | 85.22 | 75.27 | |
| DualNILM | 99.84 | 96.26 | 97.11 | 96.64 | |
| Injection Disaggregation | FHMM | Avg. RMSE: 0.8093 | Avg. MAE: 0.2200 | ||
| Seq2Seq | Avg. RMSE: 0.6834 | Avg. MAE: 0.2370 | |||
| DAE | Avg. RMSE: 0.8115 | Avg. MAE: 0.2171 | |||
| UNetNILM | Avg. RMSE: 0.7589 | Avg. MAE: 0.1983 | |||
| DualNILM | Avg. RMSE: 0.6661 | Avg. MAE: 0.1777 | |||
E.3.2. Full Results on House 2
See Table 10.
| Appliance | Method | Accuracy (%) | Recall (%) | Precision (%) | F1 (%) |
| Microwave | FHMM | 99.56 | 0.00 | 0.00 | 0.00 |
| XGBoost | 99.93 | 80.21 | 99.08 | 87.95 | |
| Seq2Point | 99.51 | 91.55 | 42.30 | 57.73 | |
| CNN-LSTM | 99.63 | 0.00 | 0.00 | 0.00 | |
| Transformer | 99.63 | 0.00 | 0.00 | 0.00 | |
| UNetNILM | 99.70 | 24.67 | 77.10 | 37.26 | |
| DualNILM | 99.91 | 79.18 | 93.96 | 85.67 | |
| Fridge | FHMM | 99.99 | 0.00 | 0.00 | 0.00 |
| XGBoost | 98.97 | 98.34 | 99.35 | 98.84 | |
| Seq2Point | 87.57 | 74.39 | 97.04 | 84.13 | |
| CNN-LSTM | 97.34 | 97.27 | 96.73 | 97.00 | |
| Transformer | 93.64 | 91.38 | 93.99 | 92.65 | |
| UNetNILM | 99.05 | 99.32 | 98.53 | 98.92 | |
| DualNILM | 99.15 | 99.94 | 98.20 | 99.06 | |
| Dish Washer | FHMM | 99.30 | 0.00 | 0.00 | 0.00 |
| XGBoost | 99.92 | 99.48 | 90.22 | 94.61 | |
| Seq2Point | 99.29 | 0.00 | 0.00 | 0.00 | |
| CNN-LSTM | 99.29 | 0.00 | 0.00 | 0.00 | |
| Transformer | 99.29 | 0.00 | 0.00 | 0.00 | |
| UNetNILM | 99.57 | 32.67 | 33.33 | 33.00 | |
| DualNILM | 99.92 | 96.73 | 94.06 | 95.13 | |
| Micro-inverter | FHMM | 1.46 | 2.31 | 1.36 | 1.71 |
| XGBoost | 94.91 | 91.83 | 93.68 | 92.73 | |
| Seq2Point | 79.33 | 45.65 | 94.42 | 61.54 | |
| Transformer | 99.13 | 99.73 | 97.92 | 98.82 | |
| UNetNILM | 99.96 | 99.99 | 99.90 | 99.95 | |
| DualNILM | 97.61 | 95.32 | 98.02 | 96.65 | |
| Avg. | FHMM | 75.08 | 0.58 | 0.34 | 0.43 |
| XGBoost | 98.43 | 92.47 | 95.58 | 93.53 | |
| Seq2Point | 91.43 | 52.90 | 58.44 | 50.85 | |
| CNN-LSTM | 98.86 | 49.19 | 48.74 | 48.96 | |
| Transformer | 97.92 | 47.78 | 47.98 | 47.87 | |
| UNetNILM | 99.57 | 64.16 | 77.22 | 67.28 | |
| DualNILM | 99.70 | 92.48 | 94.00 | 93.11 | |
| Injection Disaggregation | FHMM | Avg. RMSE: 0.8403 | Avg. MAE: 0.3367 | ||
| Seq2Seq | Avg. RMSE: 0.7217 | Avg. MAE: 0.3287 | |||
| DAE | Avg. RMSE: 0.8448 | Avg. MAE: 0.3356 | |||
| UNetNILM | Avg. RMSE: 2.2077 | Avg. MAE: 0.6601 | |||
| DualNILM | Avg. RMSE: 0.7065 | Avg. MAE: 0.2768 | |||
E.3.3. Full Results on House 3
See Table 11.
| Appliance | Method | Accuracy (%) | Recall (%) | Precision (%) | F1 (%) |
| Microwave | FHMM | 97.95 | 16.22 | 2.77 | 4.73 |
| XGBoost | 99.78 | 42.67 | 52.10 | 46.80 | |
| Seq2Point | 99.80 | 25.54 | 66.67 | 34.37 | |
| CNN-LSTM | 99.75 | 0.00 | 0.00 | 0.00 | |
| Transformer | 99.75 | 0.00 | 0.00 | 0.00 | |
| UNetNILM | 99.70 | 0.00 | 0.00 | 0.00 | |
| DualNILM | 99.88 | 65.70 | 93.19 | 75.20 | |
| Fridge | FHMM | 97.33 | 4.74 | 1.44 | 2.21 |
| XGBoost | 87.11 | 67.22 | 97.50 | 79.56 | |
| Seq2Point | 93.89 | 93.92 | 90.19 | 91.97 | |
| CNN-LSTM | 85.64 | 68.17 | 93.58 | 78.25 | |
| Transformer | 93.35 | 96.98 | 85.76 | 91.01 | |
| UNetNILM | 96.51 | 99.66 | 91.25 | 95.26 | |
| DualNILM | 98.82 | 99.00 | 98.13 | 98.56 | |
| Dish Washer | FHMM | 97.65 | 0.00 | 0.00 | 0.00 |
| XGBoost | 99.81 | 24.60 | 32.31 | 27.93 | |
| Seq2Point | 99.60 | 0.00 | 0.00 | 0.00 | |
| CNN-LSTM | 99.60 | 0.00 | 0.00 | 0.00 | |
| Transformer | 98.96 | 11.46 | 10.17 | 10.78 | |
| UNetNILM | 99.60 | 0.00 | 0.00 | 0.00 | |
| DualNILM | 99.93 | 29.00 | 31.84 | 30.35 | |
| Washing Machine | FHMM | 98.02 | 0.00 | 0.00 | 0.00 |
| XGBoost | 99.93 | 96.63 | 100.00 | 98.28 | |
| Seq2Point | 98.18 | 0.00 | 0.00 | 0.00 | |
| CNN-LSTM | 98.18 | 0.00 | 0.00 | 0.00 | |
| Transformer | 98.18 | 0.00 | 0.00 | 0.00 | |
| UNetNILM | 99.56 | 100.00 | 82.45 | 90.29 | |
| DualNILM | 99.86 | 93.51 | 99.29 | 96.25 | |
| Micro-inverter | FHMM | 28.27 | 1.67 | 0.30 | 0.50 |
| XGBoost | 87.80 | 0.00 | 0.00 | 0.00 | |
| Seq2Point | 99.02 | 94.45 | 97.66 | 96.02 | |
| CNN-LSTM | 96.40 | 76.06 | 97.27 | 84.87 | |
| Transformer | 98.79 | 92.62 | 98.55 | 95.38 | |
| UNetNILM | 99.99 | 99.90 | 99.96 | 99.93 | |
| DualNILM | 99.96 | 99.90 | 99.83 | 99.86 | |
| Avg. | FHMM | 83.84 | 4.53 | 0.90 | 1.49 |
| XGBoost | 94.89 | 46.22 | 56.38 | 50.51 | |
| Seq2Point | 98.10 | 42.78 | 50.90 | 44.47 | |
| CNN-LSTM | 95.92 | 28.85 | 38.17 | 32.62 | |
| Transformer | 97.81 | 40.21 | 38.89 | 39.43 | |
| UNetNILM | 99.07 | 59.91 | 54.73 | 57.10 | |
| DualNILM | 99.69 | 77.42 | 84.46 | 80.04 | |
| Injection Disaggregation | FHMM | Avg. RMSE: 0.5575 | Avg. MAE: 0.1766 | ||
| Seq2Seq | Avg. RMSE: 0.4974 | Avg. MAE: 0.1940 | |||
| DAE | Avg. RMSE: 0.5606 | Avg. MAE: 0.1743 | |||
| UNetNILM | Avg. RMSE: 2.2517 | Avg. MAE: 0.4708 | |||
| DualNILM | Avg. RMSE: 0.4261 | Avg. MAE: 0.1422 | |||
E.3.4. Overview on REDD
See Table 12.
| Task | House | Method | Accuracy (%) | Recall (%) | Precision (%) | F1 (%) |
| State Recognition | House 1 | FHMM | 79.52 | 1.46 | 13.47 | 2.07 |
| XGBoost | 97.90 | 74.50 | 88.74 | 79.18 | ||
| Seq2Point | 97.65 | 42.20 | 56.42 | 46.07 | ||
| CNN-LSTM | 98.67 | 46.32 | 50.75 | 47.96 | ||
| Transformer | 95.99 | 56.08 | 54.10 | 53.42 | ||
| UNetNILM | 99.05 | 76.11 | 85.22 | 75.27 | ||
| DualNILM | 99.84 | 96.26 | 97.11 | 96.64 | ||
| House 2 | FHMM | 75.08 | 0.58 | 0.34 | 0.43 | |
| XGBoost | 98.43 | 92.47 | 95.58 | 93.53 | ||
| Seq2Point | 91.43 | 52.90 | 58.44 | 50.85 | ||
| CNN-LSTM | 98.86 | 49.19 | 48.74 | 48.96 | ||
| Transformer | 97.92 | 47.78 | 47.98 | 47.87 | ||
| UNetNILM | 99.57 | 64.16 | 77.22 | 67.28 | ||
| DualNILM | 99.70 | 92.48 | 94.00 | 93.11 | ||
| House 3 | FHMM | 83.84 | 4.53 | 0.90 | 1.49 | |
| XGBoost | 94.89 | 46.22 | 56.38 | 50.51 | ||
| Seq2Point | 98.10 | 42.78 | 50.90 | 44.47 | ||
[…]